451 J of energy raises the temperature of a 10.0 g block of a metal from 21°C to 45°C. Determine the specific

heat of the metal.


1
Expert's answer
2020-11-23T06:52:59-0500

Specific heat capacity of a subtance(c)=H/mdt

Where H is the heat required to raise the temperature

m=mass of the substance

dt=change in temperature (in kelvin) which is 318-294=24K

So c=451/10×24

Specific heat capacity of metal=1.88Jg^-1K^-1